### v1

# Geometric Autoencoder Priors for Bayesian Inversion: Learn First Observe Later

Abstract
Uncertainty Quantification (UQ) is paramount for inference in engineering applications. A common inference task is to recover full-field information of physical systems from a small number of noisy observations, a usually highly ill-posed problem. Critically, engineering systems often have complicated and variable geometries prohibiting the use of standard Bayesian UQ. In this work, we introduce Geometric Autoencoders for Bayesian Inversion (GABI), a framework for learning geometry-aware generative models of physical responses that serve as highly informative geometry-conditioned priors for Bayesian inversion. Following a “learn first, observe later” paradigm, GABI distills information from large datasets of systems with varying geometries, without requiring knowledge of governing PDEs, boundary conditions, or observation processes, into a rich latent prior. At inference time, this prior is seamlessly combined with the likelihood of the specific observation process, yielding a geometry-adapted posterior distribution. Our proposed framework is architecture agnostic. A creative use of Approximate Bayesian Computation (ABC) sampling yields an efficient implementation that utilizes modern GPU hardware. We test our method on: steady-state heat over rectangular domains; Reynold-Averaged Navier-Stokes (RANS) flow around airfoils; Helmholtz resonance and source localization on 3D car bodies; RANS airflow over terrain. We find: the predictive accuracy to be comparable to deterministic supervised learning approaches in the restricted setting where supervised learning is applicable; UQ to be well calibrated and robust on challenging problems with complex geometries. The method provides a flexible geometry-aware train-once-use-anywhere foundation model which is independent of any particular observation process.
1 Introduction
Many important problems in engineering deal with spatially varying quantities that must be inferred. The task of inferring full-field information from sparse noisy measurement is a type of inverse problem (Kaipio & Somersalo, 2005); these often need to be regularized due to their inherent ill-posedness (Calvetti & Reichel, 2003). Statistically interpretable regularization can be achieved through the Bayesian paradigm (Stuart, 2010), where we combine a likelihood function characterizing data fit for a particular observation process, and a prior which characterizes assumptions about the data-generating process. Bayesian inference allows for robust and principled uncertainty quantification, where we obtain a distribution on the estimated quantity of interest given data from a particular physical system. The main challenge in enacting a Bayesian program is prior specification; often, not much can be said about the prior which often leads to diffuse and vague posteriors with relatively poor predictive accuracy when compared to deterministic methods. To alleviate this vagueness, one possible strategy is to learn a suitable prior from a dataset of related but distinct physical setups where more data may be available. This is explored in general-purpose AI applications (Shwartz-Ziv et al., 2022; Feng et al., 2023; Boys et al., 2024) and physical applications (Meng et al., 2022; Akyildiz et al., 2025; Patel & Oberai, 2021; Patel et al., 2022; Laloy et al., 2018; 2017). However, in engineering, geometry plays a crucial role in determining the dynamics arising for a physical system. Hence, for truly general purpose, practical methodology, learned priors must incorporate knowledge about the geometry of the problem at hand. But changing geometries hinder the direct learning of a prior over fields as the probability spaces involved are tied to the geometry of each individual physical system, prohibiting a direct Bayesian treatment.
In this work we propose a methodology, using graph autoencoders, to learn highly informative priors as geometry-aware generative models for use in Bayesian inversion tasks. We also show how to effectively solve the Bayesian inverse problems resulting from these priors through a pushfoward model. Importantly, the proposed scheme decouples the task of “learning” and the task of ”inferring from observations”, resulting in train-once use everywhere approach which is independent of the particular observation process for solving inverse problems, a trait not shared by related supervised learning approaches. We make the following contributions:
  • (C1)
    We propose a methodology to encode geometries and physical fields into a learned latent prior; without specifying the governing PDE or boundary conditions.
  • (C2)
    We prove that solving the Bayesian inverse problem in the learned latent representation solves the same problem as the inversion in the original space with a pushforward prior.
  • (C3)
    The proposed methodology is naturally extended to the Bayesian estimation of other important information about the observation process such as learning the observational noise.
  • (C4)
    We test the proposed methodology on: a steady-state heat problem, RANS flow around airfoils of varying geometries, damped Helmholtz resonance and source localization on car geometries. We compare our methods to supervised neural network-based methods and Graph Gaussian Processes where relevant.
  • (C5)
    We demonstrate the scalability of the method on a large terrain flow problem through a multi-GPU implementation; showing potential for training GABI foundation models.
Section 1.1 discusses the inference setup in detail; Section 1.2 goes over relevant related works; Section 2 describes the proposed methodology; Section 3 discusses the implementation of the proposed method; Section 4 tests and compares the proposed method; Appendices A, B, C, elaborate on proofs, additional numerical results and implementation details, relevant discussions on VAEs.
1.1 Setup
Let $u_{n}\in\mathcal{U}_{n}:=\mathcal{U}(\mathcal{M}_{n};\mathbb{R}^{d_{u}})$ be a (possibly vector valued) function representing the full field state of a physical system on the bounded domain $\mathcal{M}_{n}\subset\mathbb{R}^{d}$. This domain may be a manifold as we will see later. In this work, $u_{n}$ will be: heat on a surface, fluid flow in a domain, and resonance patterns on an object. Here the index $n$ refers to the different physical systems which have possibly different geometries, forcing, and boundary conditions. In numerically representing these continuous systems we obtain discretized node values $\mathbf{u}_{n}\in\mathsf{U}_{n}:=\mathsf{U}(\mathsf{M}_{n};\mathbb{R}^{d_{u}})$ which live on $\mathsf{M}_{n}=(\mathsf{V}_{n},\mathcal{E}_{n})$ where $\mathsf{V}_{n}=\{x_{1},\ldots,x_{d_{n}}\}\subset\mathcal{M}_{n}$ is a set of vertices and $\mathcal{E}_{n}$ is a set of edges connecting the vertices. As such, we represent $\mathsf{M}_{n}$ as an undirected graph describing the physical geometries on which the physics is taking place, i.e. a computational mesh. We assume to have a dataset $\mathcal{D}=\{\mathbf{u}_{n},\mathsf{M}_{n}\}_{n=1}^{N}$  [^1] A dataset with multiple full-fields per geometry $\mathsf{M}_{n}$, as $\{\{\mathbf{u}_{n,i}\}_{i=1}^{N_{n}},\mathsf{M}_{n}\}_{n=1}^{N}$, can viewed as a dataset $\{\mathbf{u}_{k},\mathsf{M}_{k}\}_{k=1}^{K}$ indexed by a dummy variable $k$ such that $k$ maps to $(i,n)$.. This dataset can be constructed from simulation data, as is the case in this paper, or from full-field data from experiment such as through particle image velocimetry, digital image correlation, magnetic resonance, electrical impedance tomography, infrared thermography etc.
First, in the proposed methodology we learn a $\mathsf{M}_{n}$-dependent prior generative model for $\mathbf{u}_{n}$. Second, using this learned highly informative prior and sparse noisy observations, $\mathbf{y}_{o}\in\mathbb{R}^{d_{o}}$  [^2] When discussing observations which are associated to a geometry not in the dataset $\mathcal{D}$ we change the index from $n$ to $o$ for clarity., of a new physical system on a new geometry, $\mathsf{M}_{o}$, we solve the Bayesian inverse problem to obtain the posterior distribution $p_{\mathbf{u}_{o}|\mathbf{y}_{o}}$. We assume to know the discretized geometry $\mathsf{M}_{o}$ associated to $\mathbf{y}_{o}$. In this paper, the observations and discretized solution field are related by:
$\mathbf{y}_{o}=\mathbf{H}_{o}\mathbf{u}_{o}+\bm{\xi}_{o},\quad\bm{\xi}_{o}\sim\mathcal{N}(\mathbf{0},\sigma^{2}{\mathbf{I}}_{o});$ (1)
$\mathbf{H}_{o}:\mathbb{R}^{d_{o}}\times\mathbb{R}^{d_{u}}\rightarrow\mathbb{R}^{o}$ is the observation operator selecting the nodes to be observed. Other observation models can be considered with ease.
Using the proposed methodology, one can encode prior beliefs through geometry-dependent full-field solution datasets. Practitioners can then construct highly-informative priors from data across varying physical geometries. This prior is independent of the observation process for any particular inference task, hence this is a learn-once apply everywhere approach.
1.2 Related Works
1.2.1 Direct Map Inversion
Many interesting approaches have been recently proposed which directly learn an amortized map from observable to solution/parameter. These can be subdivided into deterministic methods Bayesian/probabilistic approaches. Most related to the setting considered in this paper, is the work in Duthé et al. (2025). Here the authors consider a graph-ML approach to learning a deterministic map from observable $\mathbf{y}_{o}$ to the underlying vector field $\mathbf{u}_{o}$. The supervised methodology used in the comparisons in this work take inspiration from this work. Another strongly related methodology can be found in Arridge et al. (2019) although here the problem of changing geometry is not directly tackled. Other noteworthy works in this area include (Adler & Öktem, 2017; Dittmer et al., 2020) Interesting works look at probabilistic schemes for constructing probabilistic maps from observable to quantity of interest (Kaltenbach et al., 2023; Baptista et al., 2024). These works do not focus on learning priors. Other works in this area include (Adler et al., 2022; Ardizzone et al., 2018).
1.2.2 Forward and Inverse Problems on Graphs
Graphs are handled in machine learning frameworks through geometric deep learning (Bronstein et al., 2021; Wu et al., 2022; Kipf, 2016). Several important works consider inverse problems where the quantity inferred lives of a graph description of the geometry. In Garcia Trillos & Sanz-Alonso (2018) the authors prove the convergence of the graph posterior to the continuum posterior subject to some conditions. In Povala et al. (2022) the adjacency structure of a mesh is leveraged to accelerate variational inference inversion. Important works consider solving the PDE forward problem through physics informed graph machine learning(Seo & Liu, 2019; Valencia et al., 2025; Wandel et al., 2020) Used in our comparisons are linear inverse problems with Gaussian process priors, this leads to graph Gaussian process methods (Borovitskiy et al., 2021; Mostowsky et al., 2024). We find it important to mention important works in the physics informed literature, where assuming complete or partial knowledge PDEs, boundary conditions, forcing etc, one can use physics residuals to learn forward and inverse maps for PDEs (Raissi et al., 2019; Gao et al., 2022; Vadeboncoeur et al., 2023; Xu et al., 2023; Rixner & Koutsourelakis, 2021). However, in this work we focus solely on methods which do not assume knowledge of the governing PDE, boundary conditions, or forcing and are entirely data-driven in their approach.
1.2.3 Statistical Autoencoders
Related to inverse problem are latent variable representations of graph data structures. In this category the seminal work of the graph VAE (Kipf & Welling, 2016; Nazari et al., 2023). We refer the reader to Appendix C for a more elaborate discussion on the relationship between VAEs and our proposed framework. Related to this are the works (Mylonas et al., 2021; Liu et al., 2019) The proposed work centres around a class of autoencoders dubbed statistical autoencoders, of which there are many variants (Zhao et al., 2019; Kolouri et al., 2019; Turinici, 2021; Tolstikhin et al., 2018). We focus in this work on the MMD (Gretton et al., 2012) variant of the autoencoder. Many other variants could be considered. Another relevant class of autoencoders are conditional autoencoders (Sohn et al., 2015), we will make use of related ideas by conditioning encoder and decoder on geometry.
1.3 Notation
We denote by $\mathbb{P}_{z}$ a probability measure on the measurable space $(\mathcal{Z},\mathcal{S})$, and by $p_{z}$ the associated density with respect to Lebesgue measure when it exists. When $\mathbb{P},\mathbb{Q}$ are empirical measures, i.e. a mixture of Dirac measures, we will sometimes refer to these measures as the “densities” $p,q$ to lighten notation. Superscripts in parentheses denote realizations of random variables. For a function $g:\mathcal{Z}\rightarrow\mathcal{U}$, the pushforward $g_{\#}p_{z}$ means $g_{\#}\mathbb{P}_{z}(A)=\mathbb{P}_{z}(g^{-1}(A))$ for all measurable sets $A$. We note a sample $u^{(n)}\sim g_{\#}p_{z}$ is obtained as $u^{(i)}=g(z^{(i)}),z^{(i)}\sim p_{z}$.
2 Methodology
We now describe the proposed methodology in detail. In Section 2.1 we carefully layout the central result allowing for the Bayesian treatment of the inference task; that is, the equivalence between the Bayesian posterior of a pushforward prior (our generative model), and the Bayesian inverse problem in the implied latent space of the pushforward prior. Following from this result, in Section 2.2 we describe our two-step procedure: first, learn a generative model to be used as a Geometry-conditioned Bayesian prior; second, observe sparse data for the inverse problem and solve the Bayesian inverse problem in the latent space to push this posterior back out to the desired physical field living on the geometry of interest. In Section 2.3 we show how to extend our methodology to estimate the noise in the observations.
2.1 Posterior from a Pushforward Prior
We present the lemma motivating our methodology; the equivalence between the Bayesian posterior with a pushforward latent prior and the pushforward of the Bayesian posterior over the latent space.
Lemma 2.1.
Let $(\mathcal{U},\mathcal{F},\mathbb{P}_{u})$, and $(\mathcal{Z},\mathcal{S},\mathbb{P}_{z})$ be measure spaces and $g:\mathcal{Z}\rightarrow\mathcal{U}$ be $(\mathcal{S,F})$-measurable. Furthermore, let $\mathbb{P}_{u|y}$ be the Bayesian posterior on $\mathcal{U}$ with likelihood proportional to $\exp(-\Phi(u;y))$ and prior $\mathbb{P}_{u}:=g_{\,\#}\mathbb{P}_{z}$. We find
$\mathbb{P}_{u|y}=g_{\,\#}\mathbb{P}_{z|y},$
where $\mathrm{d}\mathbb{P}_{z|y}(z)=1/Z(y)\exp(-{\Phi}(g(z);y))\mathrm{d}\mathbb{P}_{z}(z)$ is the Bayesian posterior on $\mathcal{Z}$.
All proofs can be found in Appendix A. This result will be explicitly related to the geometric inference setting in Theorem 2.2, showing that with a geometry agnostic latent space, constructed via a geometry aware encoder/decoder, we can solve inverse problems with a single common latent representation relevant to any number of geometries and learn highly informative priors across geometries.
2.2 Geometric Autoencoder Priors for Bayesian Inversion (GABI)
We describe the 2 step process of learning a geometry-aware prior and how to use this for solving inverse problems.
Step 1: Learn Geometric Autoencoder Prior. Let $\mathbf{z}\in\mathsf{Z}=\mathbb{R}^{d_{z}}$ be a latent vector. We pose the $\mathsf{M}_{n}$-dependent maps
$E^{\theta}_{n}:$ $\;\mathsf{U}_{n}$ $\rightarrow\mathsf{Z},\quad$ $\text{where}\quad E_{n}^{\theta}(\mathbf{u}_{n})$ $:=E^{\theta}(\mathbf{u}_{n};\mathsf{M}_{n});$ (2a)
$D^{\psi}_{n}:$ $\;\mathsf{Z}$ $\rightarrow\mathsf{U}_{n},\quad$ $\text{where}\quad D_{n}^{\psi}(\mathbf{z})$ $:=D^{\psi}(\mathbf{z};\mathsf{M}_{n}).$ (2b)
These two maps imply a geometry conditioned autoencoder. We use $\widehat{\mathbb{E}}_{\mathcal{D}}[\,\cdot\,]$ as a shorthand for $\frac{1}{N}\sum_{(\mathbf{u}_{n},\mathsf{M}_{n})\in\mathcal{D}}[\,\cdot\,]$. Set the parameters of the auto-encoder to
$\theta^{\star},\psi^{\star}$ $=\underset{\theta,\psi}{\mathrm{arg\,min}}\;\widehat{\mathbb{E}}_{\mathcal{D}}[\|\mathbf{u}_{n}-(D_{n}^{\psi}\circ E^{\theta}_{n})(\mathbf{u}_{n})\|^{2}_{2}]+\,\mathsf{d}(p^{\theta}_{\mathbf{z}},q_{\mathbf{z}}),$ (3)
where $p^{\theta}_{\mathbf{z}}=\widehat{\mathbb{E}}_{\mathcal{D}}\,[\delta_{E^{\theta}_{n}(\mathbf{u}_{n})}]$ is the empirical distribution of encoded solutions, $q_{\mathbf{z}}=\mathcal{N}(\bf 0,\mathbf{I})$, and $\mathsf{d}:\mathcal{P}(\mathsf{Z})\times\mathcal{P}(\mathsf{Z})\rightarrow[0,\infty)$ may be any suitable divergence that is well-posed for comparing empirical measures: maximum mean discrepancy (MMD), energy distance (ED), Wasserstein, or Sliced-Wasserstein distances etc. In this work, we focus on MMD.
Step 2: Observe and Sample Posterior. For new data $\mathbf{y}_{o}$ associated to a geometry $\mathsf{M}_{o}$, we re-write the data generating model in (1) in terms of a decoded latent variable as
$\mathbf{y}_{o}=\mathbf{H}_{o}D^{\psi}_{o}(\mathbf{z})+\bm{\xi}_{o};\quad\bm{\xi}_{o}\sim\mathcal{N}(\mathbf{0},\sigma^{2}{\mathbf{I}}_{o}).$ (4)
We specify our prior distribution over $\mathbf{u}_{o}$ using the trained decoder and specify the likelihood associated to this decoded observation model
$\centering p_{\mathbf{u}_{o}}:=D^{\psi}_{o\,\#}q_{\mathbf{z}};\quad\quad p^{\psi}_{\mathbf{y}_{o}|\mathbf{z}}=\mathcal{N}(\mathbf{H}_{o}D^{\psi}_{o}(\mathbf{z}),\sigma^{2}\mathbf{I}_{o}).\@add@centering$ (5)
This likelihood along with prior $q_{\mathbf{z}}$ imply a posterior
$p^{\psi}_{\mathbf{z}|\mathbf{y}_{o}}(\mathbf{z})$ $=\frac{p^{\psi}_{\mathbf{y}_{o}|\mathbf{z}}(\mathbf{y}_{o})\,q_{\mathbf{z}}(\mathbf{z})}{p_{\mathbf{y}_{o}}(\mathbf{y}_{o})}.$ (6)
Theorem 2.2.
Assuming the relevant measurability of $D^{\psi}_{o}=D^{\psi}(\,\cdot\,;\mathsf{M}_{o})$, the posterior
$p_{\mathbf{u}_{o}|\mathbf{y}_{o}}=D^{\psi}_{o\,\#}p_{\mathbf{z}|\mathbf{y}_{o}}.$ (7)
The posterior (7) is sampled by first sampling from $p_{\mathbf{z}|\mathbf{y}_{o}}^{\psi}$ in (6) and decoding with $D^{\psi}_{o}$. Hence, by first solving the inverse problem in the latent space, we can use Theorem 2.2 to easily obtain posterior samples in the decoded solution space for a given geometry. This allows us to learn a unified prior over the solution field as a generative model over varying geometries and use it to solve Bayesian inverse problems with a highly informative learned prior on new geometries.
2.3 Observational Noise Estimation
Beyond the benefits of quantifying uncertainty about model predictions, the Bayesian paradigm offers a principled framework for estimating other aspects of the data-generating model. We highlight that such estimations are completely independent of the training of the autoencoder prior – this is not the case for direct regression methods which need to have all observation variables incorporated at training time. We consider now the common task of jointly estimating the observational noise standard deviation $\sigma$, along with the field of interest $\mathbf{u}_{n}$. This is achieved through the latent joint posterior
$p_{\mathbf{z},\sigma|\mathbf{y}_{o}}(\mathbf{z},\sigma)=\frac{p_{\mathbf{y}_{o}|\mathbf{z},\sigma}(\mathbf{y}_{o})p_{\mathbf{z},\sigma_{o}}(\mathbf{z}_{o},\sigma)}{p_{\mathbf{y}_{o}}(\mathbf{y}_{o})}.$ (8)
Corollary 2.3.
We can sample the joint posterior on the solution fields through the pushforward $p_{\mathbf{u}_{o},\sigma_{o}|\mathbf{y}_{o}}=(D^{\psi}_{o}\otimes\mathrm{Id}.)_{\,\#}\,p_{\mathbf{z},\sigma_{o}|\mathbf{y}_{o}},$ where $\mathrm{Id}.$ is the identity map.
3 Implementation
1: Specify the observation vector $\mathbf{y}_{o}$, discretized geometry $\mathsf{M}_{o}$, total sample budget $N_{s}$, number of accepted samples $N_{a}$.
2: Train $E^{\theta}$, $D^{\psi}$ through (3) on $\mathcal{D}$.
3:for [Batched/Paralelized] $i=1:N_{s}$ do
4:  Sample $\mathbf{z}^{(i)}\sim q_{\mathbf{z}}$
5:  Decode ${\mathbf{u}^{\prime}_{o}}^{(i)}=D^{\psi}(\mathbf{z}^{(i)};\mathsf{M}_{o})$
6:  Apply observation process with noise: ${\mathbf{y}_{o}^{\prime}}^{(i)}=\mathbf{H}_{o}{\mathbf{u}_{o}^{\prime}}^{(i)}+{\bm{\xi}_{o}^{\prime}}^{(i)};\quad{\bm{\xi}_{o}^{\prime}}^{(i)}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{o})$.
7:  Compute $r_{i}=\|\mathbf{y}_{o}-{\mathbf{y}^{\prime}_{o}}^{(i)}\|_{2}$.
8:return return $N_{a}$ samples of $\mathbf{u}_{n}^{(1:N_{s})}$ with least residual $r_{i}$.
Algorithm 1 Inversion with GABI-ABC
Neural Networks: We make use of meshes and graph neural networks to describe and manipulate geometric information. The proposed methodology is agnostic to any particular architecture, however we need the following characteristics: Geometry aware architecture and data-structure; Mapping to and from fixed-dimensional vectors; Non-locality. We expand on these points in B. For the heat, airfoil, and car problems we make use of a Graph Convolutional Network with interleaved nonlocal averaging layers (Lanthaler et al., 2024). For the terrain flow problem we use a GEN (Li et al., 2020) GNN architecture. We detail these architectures in Appendix B.1. We also test a Transformer on the heat problem in Appendix B.3.3.
Sampling: To realize GABI, we first train the autoencoder; then, we sample from the corresponding latent posterior as in (6) and decode the resulting latent samples to obtain samples from the posterior for a given geometry. We note two alternatives for sampling from the posterior: MCMC, and ABC sampling. We find ABC sampling to be advantageous for this task due to: (i) the massive parallelization capabilities of neural networks as opposed to the sequential nature of MCMC; (ii) the specification of a well tailored prior distribution – the prior is not unnecessarily vague leading to good overlap between the prior and posterior; (iii) the low dimensionality of the observational data – removing the need to use possibly problematic summary statistics. In Algorithm 1 we outline the implementation of the GABI-ABC variant. For MCMC sampling, we use the NUTS algorithm (Hoffman et al., 2014).
Notes on Alternative Approaches: Physics Driven Inversion. These require additional knowledge aside from access to a representative dataset and information specific to each system we wish to query such as the governing PDE, boundary conditions, forcing etc. VAEs. We refer the reader to Appendix C for an in depth discussion on the unsuitability of VAEs for the task we are solving, both in terms of the overall methodology and in terms of the autoencoder used within the proposed methodology. Supervised learning/Direct Map Inversion. One can choose to learn a supervised learning direct mapping $(\mathbf{y}_{n},\mathsf{M}_{n})\mapsto\mathbf{u}_{n}$; this approach is the main point of comparison with our method. The significant disadvantage of this approach is that the observation process (in this paper $\mathbf{y}_{n}=\mathbf{H}_{n}\mathbf{u}_{n}+\bm{\xi}_{n}$) must be known at training time and kept the same at inference time (or kept the same in a distributional sense, as explained in the comparisons). However, in practice, the observation process for different physical systems can vary enormously as we measure different quantities such as moments, local averages, summary coefficients, have a variable number of observations, or vary the type and magnitude of noise etc. In contrast to this, GABI is independent of the observation process at training time and only requires this information at test time. Thus, GABI offers a train-once-use-everywhere type foundation model, hugely advantageous for practical/industry uptake of ML models. Bayesian NNs. The problem of specifying informative priors over varying geometries persists and inherits the problems of direct maps discussed above.
4 Numerics
(a) $n=1$
(b) $n=2$
(c) $n=3$
(d) $n=4$
Figure 1: Four out of 1k geometries and solutions in dataset for the steady-state heat problem.
(a) Selected query locations
(b) marginal of $p_{\mathbf{u}_{n}|\mathbf{y}_{n}}$ at specified locations
Figure 2: (a) Four selected query locations for the sampled predictive solutions given data, the full field estimations are in Figure 6. (b) The corresponding histograms for the predictive posterior at specified query locations; the dashed black line indicates the ground truth for each histogram.
4.1 Rectangles – Steady-State Heat
In this section we study the homogenous steady-state heat equation on rectangular domains with varying width, height, and boundary conditions. Thus, for a given rectangle $\mathcal{M}_{n}$, boundary condition $h_{n}$, any $u_{n}$ satisfies the steady-state heat equation. We take $\mathcal{M}_{n}\subset\mathbb{R}^{2}$ to be $\mathcal{M}_{n}=\{x\in\mathbb{R}^{2}:0<x_{1}<l_{n},0<x_{2}<w_{n}\}$. The domain $\mathcal{M}_{n}$ is defined by its length and width $(l_{n},w_{n})$ drawn from $\texttt{uniform}((0.1,0.1),(1,1))$; we specify $h_{n}$ to be zero on the bottom and left side, and drawn from $\texttt{uniform}((0.1,0),(1,1))$ on the top and right. The solution to (9) depends on $\mathcal{M}_{n},h_{n}$. We generate 1k solutions to be in the training dataset. In Figure 1 we show example geometries and solution in the dataset. Once the model in trained on this data (the loss in shown in B.3) we select a new observation vector and geometry $\mathsf{M}_{n},\mathbf{y}_{n}$ and use ABC sampling as in Algorithm 1; we show the predictive mean and standard deviation of the sampled solution fields in Figure 6. Figure 2 shows the histogram of the predictive posterior at four query locations.
Table 1: Comparison on Heat Equation in Rectangular Domain
Method Field MAE % 1 std % 2 std Train Pred.
2D Heat Rect. Domain
GABI-ABC $(u)$ $1.58\cdot 10^{-2}\pm 1.36\cdot 10^{-2}$ 80.91% 95.59% 2.62hr 0.908s
GABI-NUTS $(u)$ $1.11\cdot 10^{-2}\pm 1.04\cdot 10^{-2}$ 66.66% 96.18% 410.31s
Direct Map $(u)$ $1.25\cdot 10^{-2}\pm 1.02\cdot 10^{-2}$ 1.47hr $0.0029s$
GP (M 1/2) $(u)$ $8.46\cdot 10^{-2}\pm 4.18\cdot 10^{-2}$ 66.36% 89.45% 0.65s
GP (M 3/2) $(u)$ $8.06\cdot 10^{-2}\pm 4.85\cdot 10^{-2}$ 65.80% 87.22% 0.64s
GP (RBF) $(u)$ $1.39\cdot 10^{-1}\pm 6.66\cdot 10^{-2}$ 14.88% 29.36% 0.64s
Table 2: Comparison on Heat Equation in Rectangular Domain – Unkown Noise
Method Field/QoI MAE % 1 std % 2 std Train Pred.
GABI-ABC ($u$) $2.09\cdot 10^{-2}\pm 1.62\cdot 10^{-2}$ 75.74% 94.76% 2.62hr 0.904s
Direct Map ($u$) $2.13\cdot 10^{-2}\pm 2.27\cdot 10^{-2}$ 1.46hr 0.0030s
GP (M 1/2) ($u$) $7.73\cdot 10^{-2}\pm 5.18\cdot 10^{-2}$ 64.87% 87.50% 0.615s
GP (M 3/2) ($u$) $7.56\cdot 10^{-2}\pm 5.56\cdot 10^{-2}$ 51.58% 76.94% 0.686s
GP (RBF) ($u$) $1.40\cdot 10^{-1}\pm 6.94\cdot 10^{-2}$ 18.03% 36.69% 0.694s
GABI-ABC ($\sigma$) $7.96\cdot 10^{-1}\pm 5.74\cdot 10^{-1}$ 42.10% 69.11% “ “
Direct Map ($\sigma$)
GP (M 1/2) ($\sigma$) $1.08\cdot 10^{~0~}\pm 1.03\cdot 10^{~0~}$ “ “
GP (M 3/2) ($\sigma$) $1.10\cdot 10^{~0~}\pm 1.47\cdot 10^{~0~}$ “ “
GP (RBF) ($\sigma$) $2.01\cdot 10^{~0~}\pm 3.61\cdot 10^{~0~}$ “ “
Comparison – Known Noise. In Table 1 we compare two sampling variants of GABI with a direct regression map as well as Gaussian Process Kriging with various kernels [^3] GABI-NUTS was tested for $10^{2}$ runs not $10^{3}$ due to prediction runtime.; M stands for Matérn. In this setup we keep number of observations at 10 randomly selected locations and the noise standard deviation at $10^{-2}$. Comparison – Unknown Noise. In Table 2 we test the noise estimation capabilities of GABI. We have the noise standard deviation be random and drawn from a shifted log-normal distribution as $\bm{\xi}_{n}=\exp(\bm{\varepsilon}-4)+10^{-3};\bm{\varepsilon_{n}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We estimate the noise for each data $\mathbf{y}_{n}$ with the GABI approach outlined in 2.3.
4.2 Airfoils – Flow Field
Table 3: Comparison on Airfoil
Method Field MAE % 1 std % 2 std Train Pred.
GABI-ABC ($p$) $6.92\cdot 10^{-2}\pm 4.39\cdot 10^{-2}$ 78.77% 97.28% 8.61hr 34.76s
Direct Map ($p$) $5.35\cdot 10^{-2}\pm 3.07\cdot 10^{-2}$ 4.68hr 0.0040s
GABI-ABC ($v_{x}$) $1.31\cdot 10^{-1}\pm 3.62\cdot 10^{-2}$ 80.36% 97.28% “ “ “ “
Direct Map ($v_{x}$) $9.30\cdot 10^{-2}\pm 1.93\cdot 10^{-2}$
GABI-ABC ($v_{y}$) $3.94\cdot 10^{-2}\pm 2.20\cdot 10^{-2}$ 75.87% 96.08% “ “ “ “
Direct Map ($v_{y}$) $3.24\cdot 10^{-2}\pm 1.12\cdot 10^{-2}$
(a) Pressure GT
(b) Pressure Mean
(c) Pressure Stddev.
(d) Pressure Error
Figure 3: Comparison of ground truth (GT), inferred mean, error, and standard deviation for pressure ($p$). The red dots correspond to the observation location for the pressure. Full results including reconstruction of vertical and horizontal velocity fields are in Figure 10
Table 4: Comparison on Car Resonance – Full Field Reconstruction and Source Localization
Method Field MAE % 1 std % 2 std Train Pred.
GABI-ABC ($u$) $2.32\cdot 10^{-1}\pm 7.21\cdot 10^{-2}$ 75.48% 96.88% 2.29hr 18.14s
Direct Map ($u$) $6.03\cdot 10^{-1}\pm 1.46\cdot 10^{-1}$ 1.22hr 0.0032s
GABI-ABC ($f$) $2.26\cdot 10^{-3}\pm 4.36\cdot 10^{-4}$ 80.55% 97.49%
Direct Map ($f$) $2.46\cdot 10^{-1}\pm 8.76\cdot 10^{-2}$ “ “ “ “
In this section we test the proposed methodology on the full field reconstruction or airflow around an airfoil from a dataset of Reynolds averaged Navier-Stokes simulations. We use the setup from Duthé et al. (2025). Here the geometric autoencoder is trained on the pressure and 2D velocity field of 1k airfoil geometries with various inflow conditions. Once trained, the observation vector $\mathbf{y}_{o}$ is taken to be noisy pressure observations at a small number of nodes on the surface of the airfoil, emulating a physical scenario when such pressure sensors may be attached to a wing/blade. The Bayesian inversion task is to reconstruct the pressure and velocity field around the airfoil, a task far too ill-posed without a highly informative prior. To compile the results in Table 3 we test GABI-ABC against the Direct Map with the same architecture type – we randomize the number of observation locations to be drawn between 5 and 50 observations with probability inversely proportional to the number of observation locations. For the supervised Direct Map approach the distribution of the number of observation must be known at train time, this is not the case for GABI who’s training is independent of the observation process. Figure 3 shows a select number of results.
4.3 Car Body – Acoustic Vibration and Source Localization
Here, we study an acoustic resonance problem with a source localization task. We have a collection of car body geometries which undergo damped vibration due to a localized Gaussian bump forcing present in the first $1/5$ of the vehicle (where the engine is), the inversion task is to reconstruct the full field acoustic vibration amplitude, $u$, as well as reconstruct the full field forcing, $f$ , on the surface of the vehicle from a small number of sparse noisy measurements of the vibration amplitude, we do not observe the forcing field. The car geometries are taken from (Umetani & Bickel, 2018). We train the model on 500 car geometries, the number of observation location is randomized as in Section 4.2. The numerical results are in Table 4 and we show a reconstruction result in 4.
(a) $u$ truth
(b) $u$ pred. mean
(c) $u$ pred. stddev
(d) $u$ pred. error
(e) $f$ truth
(f) $f$ pred. mean
(g) $f$ pred. stddev
(h) $f$ pred. error
Figure 4: Ground truth, inferred mean, stddev., and error for ampltitude $u$, and forcing $f$. In magenta are the observation locations.
4.4 Terrain – Flow Field
(a) Pressure GT
(b) Pressure Mean
(c) Pressure Std
(d) Pressure Error
(e) $\|{v}\|$ GT
(f) $\|{v}\|$ Mean
(g) $\|{v}\|$ Std
(h) $\|{v}\|$ Error
Figure 5: Ground truth, inferred mean, error, and standard deviation for pressure and the magnitude of the velocity vector, ($\|{v}\|$) across two terrains. The red dots correspond to the observation locations.
To demonstrate the scalability of our approach, we apply it to the reconstruction of flow fields over complex terrain. Here, we use a large dataset (127GB) of more than 5k RANS simulations, produced by Achermann et al. (2024). From the original $64^{3}$ grids, we extract 8-voxel-thick, terrain-conforming layers, resulting in $64\times 64\times 8$ subdomains. Each subdomain is then converted into a graph, where voxels become nodes connected to their adjacent neighbors. We scale both the training and the inference methods with multi-GPU pipelines. Using four RTX4090 GPUs, training takes 52hr while inference takes 480s. Figure 5 shows the GABI-ABC pressure and velocity results.
5 Conclusion
We propose a methodology for learning highly informative priors for solving Bayesian inverse problems across varying geometries. The scheme is realized through a geometric autoencoder which learns informative latent representations allowing for sharing of information across geometries. Once trained this foundation model can be used with any observation process to perform Bayesian full field reconstruction over new geometries. A creative use of ABC sampling leads to an efficient GPU implementation. The method achieves similar predictive accuracy to supervised direct map methods, all while outputting a full Bayesian posterior over the fields of interest and being far more flexible due to the model being independent from the observation process. The method is tested on a wide range of physical setups and geometries.
Acknowledgements
AV is supported throughthe EPSRC ROSEHIPS grant [EP/W005816/1]. MG is supported by a Royal Academy of Engineering Research Chair and EPSRC grants [EP/X037770/1, EP/Y028805/1, EP/W005816/1, EP/V056522/1, EP/V056441/1,EP/T000414/1 and EP/R034710/1]. GD and EC are supported by the BRIDGE Discovery Program of the Swiss National Science Foundation and Innosuisse (Grant No. 40B2-0_187087), as well as the French-Swiss project MISTERY funded by the French National Research Agency (ANR PRCI Grant No. 266157) and the Swiss National Science Foundation (Grant No. 200021L_212718).
Appendix A Proofs
A.1 Proof of Lemma 2.1
Proof.
Remember
$Z(y)=\int_{\mathcal{U}}\exp\left(-\Phi(u;y)\right)\mathrm{d}\mathbb{P}_{u};\quad\quad\mathbb{P}_{u}:=g_{\,\#}\mathbb{P}_{z}.$
Thus, assuming the relevant measurabilities,
$Z(y)$ $=\int_{\mathcal{U}}\exp\left(-\Phi(u;y)\right)\mathrm{d}(g_{\#}\mathbb{P}_{z})$
$=\int_{\mathcal{Z}}\exp\left(-\Phi(g(u);y)\right)\mathrm{d}\mathbb{P}_{z},$
where second equality is given by Theorem 3.6.1 in Bogachev & Ruas (2007). Applying the same equality, $\forall A\in\mathcal{F}$,
$\mathbb{P}_{u|y}(A)$ $=\frac{1}{Z(y)}\int_{A}\exp\left(-\Phi(u;y)\right)\mathrm{d}(g_{\,\#}\mathbb{P}_{z})$
$=\frac{1}{Z(y)}\int_{\mathcal{U}}\mathbbm{1}_{A}(u)\exp\left(-\Phi(u;y)\right)\mathrm{d}(g_{\,\#}\mathbb{P}_{z})$
$=\frac{1}{Z(y)}\int_{\mathcal{Z}}\mathbbm{1}_{A}(g(z))\exp\left(-\Phi(g(z);y)\right)\mathrm{d}\mathbb{P}_{z}$
$=\frac{1}{Z(y)}\int_{\mathcal{Z}}\mathbbm{1}_{g^{-1}(A)}(z)\exp\left(-\Phi(g(z);y)\right)\mathrm{d}\mathbb{P}_{z}$
$=\frac{1}{Z(y)}\int_{g^{-1}(A)}\exp\left(-\Phi(g(z);y)\right)\mathrm{d}\mathbb{P}_{z}$
$=\mathbb{P}_{z|y}(g^{-1}(A))$
$=(g_{\,\#}\mathbb{P}_{z|y})(A),$
where $\mathbbm{1}_{A}$ is the indicator function. Hence, $\mathbb{P}_{u|y}=g_{\#}\mathbb{P}_{z|y}$. ∎
A.2 Proof of Theorem 2.2
Proof.
Using Lemma 2.1, replacing $g$ with $D^{\psi}_{o}$, identifying $z,u,y$ with $\mathbf{z},\mathbf{u}_{o},\mathbf{y}_{o}$, we obtain the desired posterior. ∎
A.3 Proof of Corollary 2.3
Proof.
This follows directly from Lemma 2.1 and Theorem 2.2 by identifying $z$ with $(\mathbf{z},\sigma)$, $\mathbb{P}_{z}$ with $(\mathbb{P}_{\mathbf{z}}\otimes\mathbb{P}_{\sigma})$, $g$ with $(D^{\psi}_{n}\otimes\mathrm{Id}.)$, changing to potential function to $\Phi(u,\sigma;y)=\frac{1}{2\sigma^{2}}\|y-Hu\|^{2}_{2}$, and finally noting $\mathrm{Id}._{\#}\mathbb{P}_{\sigma}=\mathbb{P}_{\sigma}$. ∎
Appendix B Additional Numerical Results and Details
The proposed methodology is architecture agnostic, however for an architecture to be appropriate for the task we require a few particular qualities:
  • Geometry aware architecture and data-structure. As the main premise of this work is to share information about physical fields which depend on the geometry of the problem to which they relate, it is essential that the data-structure reflect this geometric information, and that the neural networks used be effective at processing this information.
  • Mapping to and from fixed-dimensional vectors. This framework relies on learning latent representations between solution to PDEs and the geometries on which these are defined. Hence, one needs to map to and from fixed dimensional latent spaces from arbitrary geometries.
  • Non-locality. All physical processes in this paper behave in non-local ways, the solution field at one point in the domain affects the solution at all other points in the domain. Hence, when encoding solution fields for a given geometry, the encoder/decoder should reflect this non-locality.
We use three architectures, a Graph Convolutional Network (explained in B.1), a Generalized Aggregation Network (explained in B.2), and a transformer (explained and tested in B.3.3).
B.1 Graph Convolutional Network
We use graph convolutional layers (Kipf, 2016) in the examples heat on rectangles, RANS around airfoil, Helmholtz on car. To induce non-locality of mappings across channels, $c$, at each layer we set $\texttt{GCN}_{l}:\mathsf{M}_{n}\times\mathbb{R}^{2c}\rightarrow\mathsf{M}_{n}\times\mathbb{R}^{c}$ and the output is concatenated with a graph level channel wise average which is expanded to populate each node of the graph. The edge attributes are the distance between nodes. The input node values always contain the node coordinates, and, for the encoder, also the data solution at the node.
In GABI: In the encoder $E^{\theta}$ we map attributed graphs to latent vectors of fixed dimensions. The decoder maps a pairing between a latent variable and a graph and populates the node attributes on the input graph. Hence the last layer in a channel wise averaging across the graph and the channel dimension is taken to be the latent dimension. The decoder $D^{\psi}$ takes as argument a graph where the node attributes are only the coordinates, hence only contains geometric information, and we expand the latent sample $\mathbf{z}$ to decode to populate all nodes with channel dimension equal the latent dimension.
In Direct Map: For the direct map method we use the same GCN layers as explained above. However now we must map observation vectors $\mathbf{y}_{n}$ and geometry $\mathsf{M}_{n}$ directly to the solution on the geometry, $\mathbf{u}_{n}$. To do this, we make the input graph node attributes the coordinates of the nodes as per the previous section, the observed nodal values $\mathbf{y}_{n}$ (where other values are 0), and a one-hot encoding indicates where this node is observed or not.
GP Regression: We use a graph Matérn Gaussian process (Mostowsky et al., 2024) with maximum marginal likelihood estimation of the hyperparameters.
The rectangle, airfoil, and car examples are run on a single RTX4090 GPU. The terrain example is run in a multi-GPU manner.
B.2 Generalized Aggregation Network (GEN)
For the graph-based autoencoder applied to the flow over complex terrain, we replace standard GCN layers with Generalized Aggregation Network (GEN) layers (Li et al., 2020). This change allows us to add more complex geometrical information to the edge features (relative coordinates, similarly to (Pfaff et al., 2020)), as GEN layers, unlike GCNs, can process multidimensional edge features. This message-passing formulation also uses a softmax-based message aggregation scheme, which allows the model to dynamically weigh the importance of messages from neighboring nodes. Importantly, we maintain the same non-local pooling operation for each layer as for the GCN-based approach.
B.3 Heat
(a) Ground Truth
(b) Predictive mean
(c) Predictive stddev
(d) Predictive Error
Figure 6: For the heat setup, we overlay in red the observation locations. (a) ground truth (b) decoded mode of predictive posterior. (c) empirical standard deviation of decoded posterior samples. (d) error between decoded mode and ground truth.
In this section, we add numerical information and results for the steady-state heat problem. The equation used to generate the dataset is
$\Delta u(x)$ $=0,\quad$ $x\in\mathcal{M}_{n},$ (9a)
$u(x)$ $=h_{n}(x),\quad$ $x\in\partial\mathcal{M}_{n}.$ (9b)
where $h_{n}$ is described in Section 4.1. We use a finite element solver over an unstructured triangular mesh to solve the systems to be in the dataset.
B.3.1 Implementation Details
Train: We use a 6-layer GCN as described in B.1 with a 100 dimensional channel space and latent space. We use 1k training geometries and train the model for 20k iterations using the Adam optimizer (Adam et al., 2014) with a learning rate of $10^{-3}$. We use a training batch size of 100.
Pred: Once trained, at inference time, in GABI-ABC we decode 10k samples in 100 batches of 100 samples. We keep 100 samples as being drawn from the posterior.
B.3.2 Additional Results
In Figure 6 we show the results of inference with GABI-ABC on for the full field reconstructing based on the small number of sparse noisy observations in red. In Figure 7 we show the two terms in the loss for training the GABI autoencoder as well as a histogram of the encoded data batch overlaid with the 1D standard normal distribution. We then test the performance of ABC sampling vs NUTS sampling as we vary the dimension of the latent space for the heat inversion task. The MAE, quality of UQ, and inference times are shown in Figure 8.
(a) Loss
(b) Latent histogram of encoded training batch
Figure 7: (a) Lose functions during training – the total loss is the sum of these. (b) histogram of the latent $p_{\mathbf{z}}^{\theta}$ across all dimensions after training.
(a) MAE
(b) UQ
(c) Inference time
Figure 8: Sampling schemes on the same autoencoder model with the same channel space and different latent dimensions. We compare (a) mean absolute predictive error, (b) uncertainty quantification under a Gaussian assumption, (c) sampling time, .
B.3.3 Transformer GABI
Table 5: Comparison of Graph Convolutional Network (GCN) and Transformer (T) architectures: Heat Equation in Rectangular Domain
Method Field MAE % 1 std % 2 std Train Pred.
T: GABI-ABC $(u)$ $1.19\cdot 10^{-2}\pm 8.65\cdot 10^{-3}$ 82.97% 97.97% 18.89hr 10.17s
T: Direct Map $(u)$ $1\cdot 10^{-2}\pm 6.69\cdot 10^{-3} 11.00$hr $0.0063s$
GCN: GABI-ABC $(u)$ $1.58\cdot 10^{-2}\pm 1.36\cdot 10^{-2}$ 80.91% 95.59% 2.62hr 0.908s
GCN: Direct Map $(u)$ $1.25\cdot 10^{-2}\pm 1.02\cdot 10^{-2}$ 1.47hr $0.0029s$
As our framework is architecture-agnostic, we also test a Transformer-based variant. Transformers can function as neural operators (Cao, 2021), achieving non-locality through attention mechanisms. Unlike the graph-based approach, this variant treats the input as an unordered point set with coordinates, without exploiting the underlying graph structure. We implement 4-layer Transformers for the encoding and decoding networks with 64-dimensional embeddings and 8 attention heads per layer, resulting in a comparable number of trainable parameters: $\sim 200$k for GABI and $\sim 100$k for the direct map. As shown in Table 5, we obtain similar results to the GCN-based approach (with the same training setup), but much slower training and prediction times, which can be attributed to the dense self-attention operations of the Transformer layers.
B.4 Airfoil
B.4.1 Implementation Details
Train: We use an 8-layer GCN as described in B.1 with a 100 dimensional channel space and latent space. We use 1k training geometries and train the model for 10k iterations using the Adam optimizer (Adam et al., 2014) with a learning rate of $10^{-3}$ with an exponential decay learning rate with decay rate 0.999. We use a training batch size of 100.
Pred: Once trained, at inference time, in GABI-ABC we decode 50k samples in 100 batches of 500 samples. We keep 100 samples as being drawn from the posterior.
B.4.2 Additional Results
In Figure 9 we show four samples from the full field dataset. In Figure 10 we show a complete example of results from the airfoil inversion task. We only measure the noisy pressure along the surface of the airfoil on the red scatter points and infer from this the full pressure field and 2D velocity field. In Figure 11 we show four random samples drawn from the generative prior model for a given geometry, demonstrating the physical coherence and diversity of the prior.
(a) $p$, $n=1$
(b) $v_{x}$, $n=1$
(c) $v_{y}$, $n=1$
(d) $p$, $n=2$
(e) $v_{x}$, $n=2$
(f) $v_{y}$, $n=2$
(g) $p$, $n=3$
(h) $v_{x}$, $n=3$
(i) $v_{y}$, $n=3$
(j) $p$, $n=4$
(k) $v_{x}$, $n=4$
(l) $v_{y}$, $n=4$
Figure 9: Four samples from the Airfoil dataset for pressure, horizontal and vertical velocity fields.
(a) Pressure GT
(b) $v_{x}$ GT
(c) $v_{y}$ GT
(d) Pressure Mean
(e) $v_{x}$ Mean
(f) $v_{y}$ Mean
(g) Pressure Std
(h) $v_{x}$ Std
(i) $v_{y}$ Std
(j) Pressure Error
(k) $v_{x}$ Error
(l) $v_{y}$ Error
Figure 10: Comparison of ground truth (GT), inferred mean, error, and standard deviation for pressure ($p$), horizontal velocity ($v_{x}$), and vertical velocity ($v_{y}$). Columns group physical variables; rows show different prediction outputs.
(a) $p$, $i=1$
(b) $v_{x}$, $i=1$
(c) $v_{y}$, $i=1$
(d) $p$, $i=2$
(e) $v_{x}$, $i=2$
(f) $v_{y}$, $i=2$
(g) $p$, $i=3$
(h) $v_{x}$, $i=3$
(i) $v_{y}$, $i=3$
(j) $p$, $i=4$
(k) $v_{x}$, $i=4$
(l) $v_{y}$, $i=4$
Figure 11: Airfoil field samples drawn from the geometry conditioned joint prior over pressure, $v_{x}$, and $v_{y}$ $(\mathbf{u}_{m}\sim D^{\psi}_{m\,\#}q_{\mathbf{z}})$ for a given geometry.
B.5 Car
For the car resonance and source localization problem we create the dataset by solving the damped Helmholtz equation (Liu et al., 2024) on a closed manifold
$\left(\Delta_{\mathcal{M}_{n}}-\kappa+i\gamma\kappa\right)u(x)$ $=f(x),$ $\quad x\in\mathcal{M}_{n}.$ (10)
B.5.1 Implementation Details
Train: We use an 6-layer GCN as described in B.1 with a 100 dimensional channel space and latent space. We use 500 training geometries and train the model for 10k iterations using the AdamW optimizer (Loshchilov & Hutter, 2017) with an initial learning rate of $10^{-3}$ with and cosine decay learning minimum rate of $10^{-5}$. We use a training batch size of 100.
Pred: Once trained, at inference time, in GABI-ABC we decode 50k samples in 100 batches of 500 samples. We keep 100 samples as being drawn from the posterior.
B.5.2 Additional Results
In Figure 12 we show some example geometries along with the forcing function to and resonance field to reconstruct. In Figure 13 we show 3 random draws from the prior for a given geometry to show the diversity of solutions and forcing fields.
(a) $u$, $n=1$
(b) $u$, $n=2$
(c) $u$, $n=3$
(d) $f$, $n=1$
(e) $f$, $n=2$
(f) $f$, $n=3$
Figure 12: Three example meshes in dataset
(a) $u$, $i=1$
(b) $u$, $i=2$
(c) $u$, $i=3$
(d) $f$, $i=1$
(e) $f$, $i=2$
(f) $f$, $i=3$
Figure 13: Given a geometry, we show 3 samples from the conditional prior.
B.6 Terrain
B.6.1 Implementation Details
Train: For both the encoding and decoding layers of the autoencoder we use an encode-process-decode backbone (Pfaff et al., 2020; Duthé et al., 2025), where each processing block is composed of a 6-layer GEN-based GNN. Both edge and node features are projected to latent vectors with 64 dimensions. We train the model for around 300 epochs using the Adam optimizer with an exponential learning rate decay (initial learning rate of $10^{-3}$, decay exponent of 0.99). This takes around 52 hours on our distributed training pipeline across four RTX4090 GPUs with an overall batch size of 20 graphs.
We process the training samples from Achermann et al. (2024), which are larger ($96\times 96\times 64$) than the testing samples ($64\times 64\times 64$). Following their procedure, we apply random on-the-fly crops and rotations during training. We then further reduce the samples by keeping only the 8 fluid cell layers directly above the terrain, resulting in a final topography-conforming training geometry of $64\times 64\times 8$.
Pred: At inference time, we decode 50k samples in batches of 40 samples, distributed over 4 GPUs, which takes around 480 seconds. We keep 100 samples as being drawn from the posterior.
B.6.2 Additional Results
In Figure 14 we show some examples of flows over different terrain geometries taken from the training set. In Figure 15 we show four random samples drawn from the generative prior model for a given terrain geometry, demonstrating the physical coherence and diversity of the prior. Due to the nature of the training data (inflows from all directions are possible), the generated samples are more diverse in nature than the previous numerical setups.
(a) $p$, $n=1$
(b) $v_{x}$, $n=1$
(c) $v_{y}$, $n=1$
(d) $v_{z}$, $n=1$
(e) $p$, $n=2$
(f) $v_{x}$, $n=2$
(g) $v_{y}$, $n=2$
(h) $v_{z}$, $n=2$
(i) $p$, $n=3$
(j) $v_{x}$, $n=3$
(k) $v_{y}$, $n=3$
(l) $v_{z}$, $n=3$
(m) $p$, $n=4$
(n) $v_{x}$, $n=4$
(o) $v_{y}$, $n=4$
(p) $v_{z}$, $n=4$
Figure 14: Four examples of flows ($p$, $v_{x}$, $v_{y}$, $v_{z})$ over complex terrain taken from the training set.
(a) $p$, $i=1$
(b) $v_{x}$, $i=1$
(c) $v_{y}$, $i=1$
(d) $v_{z}$, $i=1$
(e) $p$, $i=2$
(f) $v_{x}$, $i=2$
(g) $v_{y}$, $i=2$
(h) $v_{z}$, $i=2$
(i) $p$, $i=3$
(j) $v_{x}$, $i=3$
(k) $v_{y}$, $i=3$
(l) $v_{z}$, $i=3$
(m) $p$, $i=4$
(n) $v_{x}$, $i=4$
(o) $v_{y}$, $i=4$
(p) $v_{z}$, $i=4$
Figure 15: Flow field samples for a given test terrain drawn from the geometry conditioned joint prior over $p$, $v_{x}$, $v_{y}$, and $v_{z}$ $(\mathbf{u}_{m}\sim D^{\psi}_{m\,\#}q_{\mathbf{z}})$
Appendix C Discussion on VAEs
C.1 Alternative Variational Formulations
In this work we propose to: first train an autoencoder to learn an informative prior in the latent space, second to sample from the posterior given observational data. At first glance, an alternative approach would be to forgo this two step approach and attempt to directly learn a probabilistic geometry-conditional autoencoder in the form of a VAE
$\theta^{\star},\psi^{\star}$ $=\underset{\theta,\psi}{\mathrm{arg\,min}}\;\mathbb{E}_{\mathcal{D}_{\mathbf{y}}}\,\mathsf{KL}\left(q^{\theta}_{\mathbf{z}|\mathbf{y}_{n}}\big\|p^{\psi}_{\mathbf{z}|\mathbf{y}_{n}}\right),$ (11)
$\mathsf{KL}\left(q^{\theta}_{\mathbf{z}|\mathbf{y}_{n}}\big\|p^{\psi}_{\mathbf{z}|\mathbf{y}_{n}}\right)$ $=\log p_{\mathbf{y}_{n}}(\mathbf{y}_{n})+\mathbb{E}_{q^{\theta}_{\mathbf{z}|\mathbf{y}_{n}}}[-\log p^{\psi}_{\mathbf{y}_{n}|\mathbf{z}}(\mathbf{y}_{n})]+\mathsf{KL}\left({{q^{\theta}_{\mathbf{z}|\mathbf{y}_{n}}}\big\|q_{\mathbf{z}}}\right).$ (12)
Upon close inspection of the loss for such a model, it is apparent that information from the full solution field $\mathbf{u}_{n}$ is not incorporated in the learning, hence one cannot hope to learn the correct relationship between $\mathbf{y}_{n}$ and $\mathbf{u}_{n}$.
On could alternatively train a model of the form
$\theta^{\star},\psi^{\star}$ $=\underset{\theta,\psi}{\mathrm{arg\,min}}\;\mathbb{E}_{\mathcal{D}_{\mathbf{y}}}\,\mathsf{KL}\left(q^{\theta}_{\mathbf{u}_{n}|\mathbf{y}_{n}}\big\|p^{\psi}_{\mathbf{u}_{n}|\mathbf{y}_{n}}\right).$ (13)
However, as $\mathbf{u}_{n}$ are each associated with different geometries $\mathcal{M}_{n}$ we cannot put a common prior distribution over these in a statistically interpretable manner. Furthermore, such a model would not learn the correct relationship between $\mathbf{y}_{n}$ and $\mathbf{u}_{n}$ as observational data is ingested in an unsupervised manner; no regression loss encourages $\mathbf{y}_{n}$ to be close to $\mathbf{u}_{n}$ in any way.
There exist other models such as UQ-VAEs (Goh et al., 2022) where the decoder is replaced by the solution operator for the assumed PDE responsible for generating the data. This of course requires us to know the governing of the problem, the boundary conditions etc – and has the same challenges and limitations as other physics-informed methodologies. Furthermore, this framework is not made to handle variable geometries.
C.2 VAEs instead of Regularized Autoencoders
One might also wonder why the proposed methodology favors the use of a regularized auto-encoder rather then a Graph-VAE. By looking at (11) (substituting $\mathbf{y}_{n}$ for $\mathbf{u}_{n}$ as in our use the AE) it is plain to see the that latent space regularization $\mathsf{KL}$ term encourages the per-sample embedding variational approximation to be close to a standard normal. In contrast to this, the statistical AE loss (3) encourages the embedded data to be close to a standard normal at a distributional level. Because our framework uses as Bayesian prior the autoencoder embedding distribution, assumed to be a standard normal, the AE approach is preferable. The Gaussianity of the embedded data distribution is explicitly encouraged in (3). Furthermore, the encoder decoder in (3) is deterministic, drastically simplifying the subsequent form of the likelihood in (5) while maintaining the probabilistic aspect of the method through Bayes’ theorem.